Accessibility Assistance

Skip to Content




Information Technology
George A. Smathers Libraries
University of Florida
Suite 1250
2124 NE Waldo Rd
Gainesville, FL 32609

Telephone: (352) 273-2800
Fax: (352) 392-2354

Information Technology Staff Meeting Notes
September 14, 2010

Present: Shruti Bagri, Thomas Bielicke, Laurie Brennan, Will Chaney, Logan Clapp, Matthew Daley, Debra Fetzer, Winston Harris, Michael Jay, Chris Nicolich, Amy Polk, Patricia Ruwell, Josh Spurgin, Evan Wack

Guest: Rachel Schipper
Announcements
  1. Open Enrollment: September 27-October 22.
  2. Eastside Open House: Friday, October 1st. The event will be announced, there will be food and we'll be open all day.
  3. TSS - topic of filling vacant positions was discussed. Judy makes the final decisions. Market Equity for faculty is overdue. The only pot of money available is carry forward from vacancies. That money has to cover a lot. Brian Keith is working on phasing in Market Equity. We'll let you know when more decisions have been made. Carry-forward from salary is recurring. Could do both, but would not be able to fill all positions. TSS - $10,000 fund - can petition to use it for something to make your life easier. Need to have templates and cost proposals out by mid-October. This fund comes from overdue book fines. Thorough justification needed.
  4. Rachel will be meeting with the new UF VP and CIO, Elias G. Eldayrie. Are there any questions we would like her to ask on our behalf? We would like to know if he plans on continuing the merge/centralization of all UF IT. What is his long-range plan for UF IT? Rachel would like to know his intent in regards to Open Source clients (Sakai, OLE). Will there be more training opportunities offered on campus?
  5. Happy Birthday, Mark Sullivan.
Review Minutes of 8/31/10
  1. Michael asks that the SW - Pending Outside Forces item (Acq Order Tracking) be removed from the Agenda.
Ongoing IT Updates

Hardware
  1. Public printing - software upgrade pushed back two weeks. Mid-October is now the target. Pharos kept pushing the date back for software tweaks. Most discussions have been via conference calls, but there are meeting minutes available.
  2. Public laptops - nothing new. Public Service folks are discussing. Fines are being re-evaluated.
  3. Staff computer replacement - started today, hope to be done by Friday.
  4. Hardware SharePoint site - we are still populating it. It's a work in progress. We want to hook it into the SQL server, but the Web part is locked down. Evan is in communication with the UFAD SharePoint guy. We hope to start making Camtasia tutorials as soon as things die down a bit around here. We've reached out to Matthew for a template design.
  5. Win7 bug - Microsoft has Evan trying some things. They suggested we sysprep the base, not just the image.
  6. Surplus property - inner vault is now empty.
  7. Print queue renaming is done.
  8. Cannon meter reads - Cindy McLean requested print reports for the last four months. Talk to Grace about it.
  9. Staff network printers - Josh is redoing the inventory in Grover.
  10. Video conferencing - the company will be out this week to work on the 1A display and tweak the Eastside unit. HVAC - AC noise - is it back? Unit behind the podium - Judy would like it hidden. Evan to look into it and get an estimate.
  11. SAN expansion - only 200G left. Need to expand the storage (at a monthly cost). Otherwise, we will have to ask folks to weed out items in their storage or we'll run out of space.
  12. Help Desk - hopefully our newly hired student, Diana Etienne, will be starting this week. Joe in HR is submitting a requisition for the job in the remaining time bracket. It has been unusually difficult to get students in here to interview.
Software
  1. Document Disposition database - working on changes with the function of panels and nightly run batch jobs. Other ASERL dean changes not yet decided.
  2. Macros - quite a few outstanding. New one involves an ALEPH cut-and-paste function.
  3. Grover mods - done.
  4. New ticket today - Laurie Taylor was unable to submit a Grover ticket using an Ipad. Chris is getting a handle on the scope of changes to make it functional.
  5. OLE - Maureen Kelly is in Chicago attending OLE meetings. She will be attending Rice training next week.
  6. New request - OCR automation - production jobs for Prime when it switches to Abby.
Network/Server Administration
  1. EPO server - Logan got an email from Trish about space on the server. He attempted to increase the drive size. The change didn't take, when he moved it back and rebooted, the server did not come back up. The data is still there. Could move it to another server and then link it up to its database. Evan would prefer it be restored and then moved. Virus protection is still working. Needs to be fixed sooner rather than later. SUS is fine for now.
  2. SharePoint - Amy met with Acquisitions & Cataloging about their sites. They basically just need blog and wiki functionality. New Horizons training - will be offered in November. There does not seem to be training for End Users, but we could get the trainers trained so they could hold classes for end users. We would prefer if folks could manage their own sites.
  3. 2nd Life - ongoing, tested. It is working better in the server environment. Hope to fix it by next week.
  4. Server virtualization - only a few left. Web server - will try to move at the end of the semester. Apps server - will move whenever Winston finds it convenient.
  5. Physical server move - Logan is going to campus tomorrow to get the East switch cleared up. Then he will go to MSL and move the server. Need to push VIVO off. Worst case scenario, will move the server from MSL to Library East. Rack to move to Eastside.
  6. DLC phone moves - due to GIG switch, the DLC phones had to move to the port on the old switch. New switch to be ready in about a week.
ILS
  1. Aleph report tickets - the Baldwin report needs to better defined. The rest are ongoing: slow but steady.
DLC General Issues
  1. OCR2 is functioning correctly. OCR4 is functioning correctly, given the hardware issues. Abby is working correctly. All working correctly within the parameters of what we have here.
  2. UFDC server configuration - Logan and staff to go over documentation with DLC on Friday. Will go over final notes on restoring from tape.
Web Projects
  1. Matthew presented the Oil Spill site. Tom Frasier liked it, but has not sent any content. He also wanted to minimize the Library branding. Matthew explained he would have to discuss that with Judy. They are trying to get the domain: oilspill.ufl.edu. They are trying to get Matthew to find the content for the site. It should be referred to Vernon for site content.
  2. Advance booking - FCLA sent Matthew a version that works. Will use it to reserve rooms.
  3. Link naming conventions - Tom is working on it. Rachel thinks we need widespread understanding. Need to make it palatable. Will notes that the Library web site is different from DLC. Matthew will get a copy of the document to Rachel.